OpenStudy (anonymous):

Use the properties of exponents to simplify each of the following expressions, writing your answer with only positive exponents.

OpenStudy (anonymous):

[100^0(x^-1y^3)^-1/x^2y]^-3

OpenStudy (anonymous):

you know what 100^0=, right?

OpenStudy (anonymous):

yes 1 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can distribute the rest of the exponents to inside the parentheses, leaving negative answers until the very end

OpenStudy (anonymous):

when you raise a power to a power, you multiply those 2 powers

OpenStudy (anonymous):

for example, (x^-1)^-1 = x^(-1*-1)=x

OpenStudy (anonymous):

now what do i do? sorry i am confused.